Operation Summary for January 8, 2010 - January 11, 2010 |
The Integrated Luminosity for the period from 1/4/10 to 1/11/10 was 47.74 pb-1. Friday Jan. 8 The day shift began with Operations monitoring the stash (266.6E10), with the Antiproton Source stacking (192.6E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments in access. During the day shift, the CDF and D0 collision hall accesses ended around 10:30 AM. Tevatron (TeV) personnel conducted studies. Operations put the TeV into shot setup at 2:15 PM and established store 7491 at 3:30 PM with an average initial luminosity (AIL) of 290E30. During the evening shift, Pbar stacked and transferred antiprotons to the Recycler as appropriate. Operations sent beam to NUMI and MiniBooNE when it was available. Saturday Jan. 9 The midnight shift began with Operations monitoring store 7491 (87.55E30) and the stash (250.5E10), with Pbar stacking (22.5E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7491 at 7:17 AM and put the TeV into shot setup. Near the end of the shift, a Pbar Accumulator kicker (A:IKIKP1) tripped off and would not reset. Operators called in Support technicians to deal with it. During the day shift, Operations established store 7492 at 8:45 AM with an AIL of 282.79E30. Support personnel had replaced the Accumulator kicker power supply by 10:15 AM, but later more problems appeared, so they moved the kicker back to operate from the in-rack power supply for A:IKIK; they did this while there was a Pbar access to fix a vacuum problem in the Pbar Debuncher D30 area. During the evening shift, Support personnel came out of Pbar around 4:05 PM and stacking resumed by 4:15 PM. Operations reported finding the I- Source ion gauge off; when it was turned on, the source output dropped; this source was in the test stand. Sunday Jan. 10 The midnight shift began with Operations monitoring store 7492 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, a TeV RF station (TRF4) tripped off on an interlock LCW flow fault, resulting in a slight loss of luminosity. Operators turned off the heaters and the rest of the station and informed RF personnel. There was some downtime due to a vacuum burst in the MIRF sector. Operations terminated store 7492 at 7:45 AM and put the TeV into shot setup. During the day shift, Operations established store 7493 at 9:30 AM with an AIL of 247.35E30. TRF4, which had been turned back on, tripped off during scraping and caused the low initial luminosity. Operators informed TeV RF technicians. During the evening shift, RF techs came in to look at TRF4. The H- Source output trailed off (from 52mA to 37mA) during the shift. Operations tuned Linac and Booster to operate more efficiently at the lower source output. Monday Jan. 11 The midnight shift began with Operations monitoring store 7493 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, operators continued to tune Linac and Booster. Operations terminated store 7493 at 12:30 AM and then put the TeV into shot setup. Operations established store 7494 at 2:06 AM with an AIL of 253.2E30. There were some issues late in the shift with Pelletron trips. Operations consulted with the Electron Cooling personnel and then made necessary adjustments to the system. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster personnel to make tune measurements. She will also allow D0 and CDF to make an access at the end of store 7494, around 3 PM. |
|
More Information Comments and Suggestions |
| last modified 02/03/2010 email Fermilab |
| Security, Privacy, Legal | |